# EvaSurf: Efficient View-Aware Implicit Textured Surface Reconstruction on Mobile Devices

### v3

@@ IV-E Additional Results
Comparison with Gaussian-based methods. Gaussian-based methods [27, 29, 28] cannot render on mobile devices due to requirement of large disk storage and lack of mobile-end optimization. Besides, the reconstructed geometry is not satisfied for further application. 2DGS [26] can reconstruct the corresponding geometry but is still unable to render on mobile devices. In contrast, EvaSurf can achieve real-time rendering on mobile devices while reconstructing accurate geometry for further applications. We compare the reconstructed geometry of our method and 2DGS [26] in Figure 9 and Table I. It can be seen that 2DGS cannot handle objects with reflective surfaces while EvaSurf reconstruct highly accurate geometry and achieve better results for all the synthetic scenes.

### v4

@@ IV-E Additional Results
Comparison with Gaussian-based methods. Gaussian-based methods [27, 29, 28, 57] cannot render on mobile devices due to requirement of large disk storage and lack of mobile-end optimization. Besides, the reconstructed geometry is not satisfied for further application. 2DGS [26] can reconstruct the corresponding geometry but is still unable to render on mobile devices directly. In contrast, EvaSurf can achieve real-time rendering on mobile devices while reconstructing accurate geometry for further applications. We compare the reconstructed geometry of our method and 2DGS [26] in Figure 9 and Table I. It can be seen that 2DGS cannot handle objects with reflective surfaces while EvaSurf reconstruct highly accurate geometry and achieve better results for all the synthetic scenes.
@@
Rendering Memory Footprint Comparison. We further provide the comparison of rendering memory footprint of different methods to demonstrate the applicability of EvaSurf. We use the NeRF synthetic dataset as the example and the results are shown in Table X. It is noteworthy that we implement a three.js based renderer for 2DGS [26] for mobile rendering. It can be seen that our method requires much less memory footprint when rendering at interactive frame rates. This can be beneficial to extending the applications of 3D reconstruction models in real-life scenarios.
TABLE X: Comparison of rendering memory footprint. We present the rendering memory footprint of different methods.
Method Memory Footprint ($\downarrow$) Rendering Speed ($\uparrow$)
Re-ReND [30] 128 MB 54 FPS
MobileNeRF [25] 199 MB 42 FPS
2DGS [26] 82 MB 50 FPS
NeRF2Mesh [31] 75 MB 52 FPS
Ours 48 MB 50 FPS
@@ VI Limitation
VII Acknowledgements
This work was supported in part by NSFC (62201342), and Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102). Authors would like to appreciate the Student Innovation Center of SJTU for providing GPUs.